OpenStudy (anonymous):

John has 15 more music CDs than Michael. Together they have a total of 55 CDs. How many CDs does Michael have?

OpenStudy (anonymous):

let M have x CDs J=x+15 M+J=55 x+15 +x =55 2x=40 x=20